OpenStudy (anonymous):

Create an expression to find the time for the given distance and rate, and find the value of the expression. The time is minutes.

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

@Qwertty123

OpenStudy (qwertty123):

ok, so you need to find the time is takes (in seconds) to go a distance of 90 miles, if you are travelling 15 mi/hr?

OpenStudy (qwertty123):

Use the speed formula: v=d/t where v is your speed, d is the distance travelled and t is the time. So, d=90 and v=15. We solve for t, keeping in mind that it will be in hours, not seconds.

OpenStudy (qwertty123):

we have, t=d/v t=90/15=6 hours. Now we need to convert this to seconds....

OpenStudy (qwertty123):

There are 60 minutes in 1 hour and 60 seconds in one minute. There are 60*60=3600 seconds in one hour. Therefore, in 6 hours, we have: 6∗3600=?